OpenStudy (espex):

You slide a box of books at constant speed up a 23° ramp, applying a force of 180 N directed up the slope. The coefficient of sliding friction is 0.18. (a) How much work have you done when the box has risen 1 m vertically? _________J (b) What is the mass of the box? _________kg

OpenStudy (espex):

I figure out that the work was 461.14J, but I am having difficulty relating it to mass.

OpenStudy (anonymous):

\[\Sigma F_x = 0\] \[F = \mu_kN\] \[\Sigma F_y=0\] \[N = mg\]

OpenStudy (anonymous):

\[m = \frac{ F }{ \mu_kg }\]

OpenStudy (espex):

Thanks @Ankh, you got me pointed in the right direction. Since it's not accelerating, the \(\sum F_x = 0\) so \(\sum F_{-x} = \sum F_{+x}\) \[\rightarrow \mu mg\cos\theta + mg\sin\theta = 180\] \[\rightarrow m(\mu g\cos\theta + g\sin\theta) = 180\] \[\rightarrow m = \frac{180}{(\mu g\cos\theta + g\sin\theta)}\] \[\rightarrow m = \frac{180}{((.18\times 9.81\times\cos(0.401)) + (9.81\times\sin(0.401)))}\] \[m = 32.997kg\]